OpenStudy (anonymous):

find the exact value of the given expression sin(5pi/3-5pi/4)

OpenStudy (ddcamp):

sin(u-v) = sin(u)cos(v)-cos(u)sin(v) u = 5pi/3 v = 5pi/4

OpenStudy (anonymous):

i did that and got radical 6 -radical 2/4

OpenStudy (ddcamp):

Hmm... I got \[\frac{ \sqrt{6}+\sqrt{2} }{ 4 }\]

OpenStudy (anonymous):

but that was marked wrong?

OpenStudy (ddcamp):

I don't know then. Unless I'm totally forgetting the unit circle, that's right.

OpenStudy (anonymous):

okay what about sin(arcsinx+arccosx) how do i simplify that?

OpenStudy (ddcamp):

sin(u+v) = sin(u)cos(v)+cos(u)sin(v) sin(asin(x)+acos(x)) sin(asin(x))cos(acos(x)) + cos(asin(x))sin(acos(x)) x² + (1 - x²) 1 sin(acos(x)) = cos(asin(x)) = sqrt(1-x²)

OpenStudy (ddcamp):

So it simplifies to 1